Find two numbers whose sum is 24 such that the sum of the square of one plus six times the other is a minimum.

1
Expert's answer
2022-05-18T10:07:38-0400

Let "x=" the first number.Then the second number will be "24-x."

Consider the function


"f(x)=x^2+6(24-x)"

Find the first derivative with respect to "x"


"f'(x)=(x^2+6(24-x))'=2x-6"

Find the critical number(s)


"f'(x)=0=>2x-6=0=>x=3"

If "x<3,f'(x)<0, f(x)" decreases.

If "x>3,f'(x)>0, f(x)" increases.

The function "f(x)" has a local minimum at "x=3."

Since the function "f(x)" has the only extremum, then the function "f(x)" has the absolute minimum for "x\in \R" at "x=3."

The first number is "3." The second number is "21."